Calculated using the value
g = 9.8 m/s2
,the time interval required for an object released from rest at some arbitrary point A,above the ground to reach the ground is Δt.

1.If g were 5.9 m/s2 , by what factor would the time interval Δt′
calculated for this value of g
differ from the value calculated using g
g = 9.8 m/s2
1
Expert's answer
2020-02-26T10:07:19-0500

Free fall is a type of uniformly accelerated motion. The formula for the distance traveled by the body in free fall released from rest can be written as

(1) "h=\frac{g\cdot (\Delta t)^2}{2}" .

Let point A be at the height h from the ground. Then we can solve the equation (1) regarding "\Delta t".

(2) "\Delta t=\sqrt{\frac{2h}{g}}"

If free fall occurs on any other planet with "g=g_{'}" we have for time interval

(3)"\Delta t^{'}=\sqrt{\frac{2h}{g_{'}}}"

Thus the time interval Δt′ differ from the value "\Delta t" calculated using g by factor

"\frac {\Delta t^{'}}{\Delta t}=\sqrt{\frac{g}{g^{'}}}=\sqrt{\frac{9.8}{5.9}}=1.3"

Answer: The time interval Δt′ for value of "g^{'}=5.9 ms^{-2}" differ from the value "\Delta t" calculated using "g=9.8ms^{-2}" by the factor 1.3. The time of free fall increased as the acceleration decrease.